# Discovering new robust local search algorithms with neuro-evolution

### v1

@@ Abstract
This paper explores a novel approach aimed at overcoming existing challenges in the realm of local search algorithms. Our aim is to improve the decision process that takes place within a local search algorithm so as to make the best possible transitions in the neighborhood at each iteration. To improve this process, we propose to use a neural network that has the same input information as conventional local search algorithms. In this paper, which is an extension of the work goudet2024emergence presented at EvoCOP2024, we investigate different ways of representing this information so as to make the algorithm as efficient as possible but also robust to monotonic transformations of the problem objective function. To assess the efficiency of this approach, we develop an experimental setup centered around NK landscape problems, offering the flexibility to adjust problem size and ruggedness. This approach offers a promising avenue for the emergence of new local search algorithms and the improvement of their problem-solving capabilities for black-box problems.
@@ 1 Introduction
Fitness landscape analysis malan2021 provides the optimization and evolutionary computation community with theoretical and practical tools to examine search landscapes. It allows for the assessment of problem characteristics and the evaluation of the performance of search algorithms. In the context of combinatorial fitness landscapes, these are represented as graphs defined by a discrete search space and a neighborhood relation. A fundamental challenge lies in developing a search algorithm capable of navigating a fitness landscape to reach the highest possible fitness value.
@@
In various works, different machine learning approaches have been used either to consider offline adjustment, selection of parameters, or online control of the search process using reinforcement learning (RL) techniques (see MamaghanMMKT22 for a recent survey). The use of neural networks (NNs) in solving combinatorial optimization problems has been studied for decades TrafalisK09 , starting with the early work of Hopfield Hopfield82 . Recent applications of Graph Neural Networks in the context of combinatorial optimization have been proposed to reach optimal solutions or to assist the solving algorithm in proving the optimality of a given solution (see CappartCK00V23 for a recent survey).
@@
In the traveling salesman problem, a GNN can be used to predict the regret associated with adding each edge to the solution to improve the computation of the fitness function of the LS algorithm HudsonLMP22 . Note that NNs are classically used in surrogate model-based optimization WillmesBJS03 . In SantanaLV23 , the authors introduce a GNN into a hybrid genetic search process to solve the vehicle routing problem. The GNN is used to predict the efficiency of search operators and to select them optimally in the solving process. A deep Q-learning approach has recently been proposed to manage the different stages of an LS-based metaheuristic to solve routing and job-shop problems FalknerTBS22 . High-level solving policies can often be managed by reinforcement learning in LS processes VeerapenHS13 . In this paper, our purpose is different, as we focus on building simple search heuristics for black-box problems, rather than scheduling specific operators or parameters. In particular, we assume that the learning process cannot be based on the immediate rewards that are used in RL. This motivates our choice of neuro-evolution. Note that this type of approach has recently been used to discover new genetic and evolutionary algorithms to solve continuous black-box optimization problems lange2023discoveringGen ; lange2023discoveringEvo . In this work, we propose to use it to discover new local search algorithms for combinatorial optimization problems.
Objective of the paper This study explores the potential for emerging search strategies to overcome existing challenges. The objective is to change how information is leveraged while retaining simple and generic search components. Considering a basic hill climber algorithm to achieve a baseline search process for solving black-box binary problem instances, our aim is to benefit from machine learning techniques to get new local search heuristics that will be built from basic search information instead of choosing a priori a predefined move heuristic (e.g., always select the best neighbor).
@@
Hence, our goal is to provide a NN with the same information as a basic LS and, after training, to use the NN as the basic move component of a simple LS process. In this work, we study different ways of representing this information, from the most basic possible representation to a more complex one that exploits an ordering relation in the neighborhood. To evaluate the efficiency of our approach, we define an experimental setup based on NK landscape problems, which allows us to describe a fitness landscape whose problem size and ruggedness (determining the number of local minima) can be adjusted as parameters. Finally, to check the robustness and generalizability of our learned strategies, we propose to evaluate them on new instances of another type of pseudo-Boolean problem, namely the quadratic unconstrained binary optimization (QUBO) problem.
@@ 2.2.2 Observations
In this work, we propose to study different observation functions that will be used by local search strategies to make decisions. Given $s\in\mathcal{S}$, with $s=(x,\mathcal{N}(x))$, with $x$ the current solution encountered by the local search and $\mathcal{N}(x)=\{flip_{i}(x)\mid i\in\llbracket 1,N\rrbracket\}$ its one-flip neighborhood, we define four observation functions from $\mathcal{S}$ to $\mathbb{R}^{N\times d}$:
@@
    The first observation function abstracts the sate as the variation of the fitness function for each possible flip of the values of $x$.
    $o_{1}(x)=\begin{pmatrix}\Delta_{1}(x)\\ \Delta_{2}(x)\\ \vdots\\ \Delta_{N}(x)\\ \end{pmatrix},$
@@
    $o_{2}(x)=\begin{pmatrix}f_{{\rm obj}}(x)&f_{{\rm obj}}(flip_{1}(x))\\ f_{{\rm obj}}(x)&f_{{\rm obj}}(flip_{2}(x))\\ \vdots&\vdots\\ f_{{\rm obj}}(x)&f_{{\rm obj}}(flip_{N}(x))\\ \end{pmatrix}$
@@
    The third observation function $o_{3}:\mathcal{S}\rightarrow\mathbb{R}^{N}$ is based on a ranking of the solutions associated with positive, negative of null variations of the fitness function for each possible flip of the values of $x$. We define $N^{+}(x)$ and $N^{-}(x)$ the numbers of solutions in $\mathcal{N}(x)$ with respectively strictly greater and lower fitness than $x$.
    Then $o_{3}(x)$ is vector of size $N$, such that for $i=1,\dots,N$:
    • $o_{3}(x)_{i}=rank^{+}(\Delta_{i}(x))/N^{+}(x)$, if $\Delta_{i}(x)>0$;
    • $o_{3}(x)_{i}=rank^{-}(\Delta_{i}(x))/N^{-}(x)$, if $\Delta_{i}(x)<0$;
    • $o_{3}(x)_{i}=0$, if $\Delta_{i}(x)=0$.
@@
    Example 2 Let us consider the following situation with $x\in S$ and $N=6$ : $\Delta_{1}(x)=1,\Delta_{2}(x)=4,\Delta_{3}(x)=-2,\Delta_{4}(x)=-5,\Delta_{5}(x)=0,\Delta_{6}(x)=-7$. We have thus $rank^{+}(\Delta_{1}(x))=1,rank^{+}(\Delta_{2}(x))=2,rank^{-}(\Delta_{3}(x))=-1,rank^{-}(\Delta_{4}(x))=-2,rank^{-}(\Delta_{6}(x))=-3$. Hence, the corresponding observations are $o_{3}(\Delta_{1}(x))=\frac{1}{2},o_{3}(\Delta_{2}(x))=1,o_{3}(\Delta_{3}(x))=-\frac{1}{3},o_{3}(\Delta_{4}(x))=-\frac{2}{3},o_{3}(\Delta_{5}(x))=0,o_{3}(\Delta_{6}(x))=-1$.
    This rank transformation always returns values in the range $[-1,1]$ and keeps the information that a flip move can deteriorate or improve the current fitness solution. An interesting property of this transformation, is that it is independent of a change in the scale of the objective function. $o_{3}(x)$ remains unchanged even if $f_{{\rm obj}}$ is multiplied by a strictly positive real value $\lambda$.
@@
    The fourth observation function $o_{4}:\mathcal{S}\rightarrow\mathbb{R}^{2N}$ uses the same ranking transformation as $o_{3}$, in addition to a z-score transformation of the $\Delta_{i}(x)$ values. We define the z-score of $\Delta_{i}(x)$ as $Z(\Delta_{i}(x))=\frac{\Delta_{i}(x)-\mu(x)}{\sigma(x)}$, with $\mu(x)$ and $\sigma(x)$ are the mean and standard deviation, respectively, of the set of $\Delta_{i}(x)$ values for $i=1,\dots,N$. This z-score indicates how many standard deviations a given fitness variation is above or below the mean of possible fitness variations. Thus,
    $o_{4}(x)=\begin{pmatrix}o_{3}(x)_{1}&Z(\Delta_{1}(x))\\ o_{3}(x)_{2}&Z(\Delta_{2}(x))\\ \vdots&\vdots\\ o_{3}(x)_{N}&Z(\Delta_{N}(x)).\\ \end{pmatrix}$
@@ 2.2.3 Deterministic policy and actions
Following a reinforcement learning-based description, a local search process can be encoded by a policy $\pi:\mathbb{R}^{N\times d}\to{\cal A}$ where ${\cal A}$ is a set of actions. Here we consider deterministic policies, i.e. for each $o\in\mathbb{R}^{N\times d}$, there exists one and only one action $a\in{\cal A}$, such that $\pi(o)=a$. Note that the policy can be parameterized by a parameter vector $\theta$. The set of all policies is $\Pi=\{\pi_{\theta}\mid\theta\in\Theta\}$ where $\Theta$ is the parameter space.
@@ 2.2.4 Transitions and trajectory
Example 1. In order to illustrate our framework, let us consider a basic hill climber (HC) that uses a simple best-improve strategy using objective function $f_{{\rm obj}}$ as fitness function. In the following, we only consider LS processes that do not involve memory that records past decisions. Hence, a state of a HC is fully described by a current solution $x\in{\cal X}$ and its neighborhood $\mathcal{N}(x)$. An observation will abstract the state as the variation of the fitness function for each possible flip of the values of $x$, $o_{2}(x)=(\Delta_{1}(x),\dots,\Delta_{N}(x))\in\mathbb{R}^{N}$. In a best-improve HC process, the search stops when a local optimum is reached and no improving move can be performed. We consider ${\cal A}=\{flip_{i}\mid i\in\llbracket 1,N\rrbracket\}$. Then, we define the policy $\pi_{HC}(o_{1}(x))=\operatorname*{argmax}_{a\in{\cal A}}(f_{{\rm obj}}(\delta(x,a))-f_{{\rm obj}}(x))$. Note that when a local optimum is reached, the search gets stuck.
@@ 2.3 Policy Learning for a Set of Instances
Figure 1: Global View of the Process
@@ 3.1 Neural network local search policy
We introduce a deterministic LS policy $\pi_{\theta}:\mathbb{R}^{N\times d}\rightarrow\mathcal{A}$, called Neuro-LS, which uses a neural network $g_{\theta}$, parametrized by a vector of real numbers $\theta$. When in a state $x$, this neural network takes as input an observation matrices $o(x)\in\mathbb{R}^{N\times d}$ and gives as output a vector $g_{\theta}(o)$ of size $N$ whose component $g_{\theta}(o)_{i}$ corresponds to a preference score associated to each observation $o_{i}$. Then, the action $a_{i}$ corresponding to the highest score $g_{\theta}(o)_{i}$ is selected.
@@
A desirable property of this neural network policy is to be permutation equivariant with respect to the input vector of observations, which is a property generally entailed by a local search algorithm, in order to make its behavior consistent for solving any type of instance. Formally, an LS algorithm is said to be invariant to permutations in the observations if for any permutation $\sigma$ on $\llbracket 1,N\rrbracket$, we have $a_{\sigma(i)}=\pi_{\theta}(o_{\sigma(1)},o_{\sigma(2)},\dots,o_{\sigma(N)})$. As an example, the basic hill climber HC defined with Example 1 in Section 2.2 has this property.
@@
where $L_{h}:\mathbb{R}^{n_{h-1}}\rightarrow\mathbb{R}^{n_{h}}$ is an affine linear map defined by $L_{h}(\mathbf{x})=\mathbf{W}_{h}\cdot\mathbf{x}+\mathbf{b}_{h}$ for given $n_{h}\times n_{h-1}$ dimensional weight matrix $\mathbf{W}_{h}$ (with coefficients $\{w^{h}_{k,l}\}_{1\leq k\leq n_{h}\atop 1\leq l\leq n_{h-1}}$), $n_{h}$ dimensional bias vector $\mathbf{b}_{h}$ (with coefficients $\{b^{h}_{k}\}_{1\leq k\leq n_{h}}$) and $\sigma:\mathbb{R}^{n_{h}}\rightarrow]-1,1[^{n_{h}}$ the element-wise nonlinear activation map defined by $\sigma(\mathbf{z}):=(tanh(z_{1}),...,tanh(z_{n_{h}}))^{T}$. We denote by $\theta$, the set of all weight matrices and bias vector of the neural network $g_{\theta}$ : $\theta:=\{(\mathbf{W}_{1},\mathbf{b}_{1}),(\mathbf{W}_{2},\mathbf{b}_{2}),\dots,(\mathbf{W}_{{H+1}},\mathbf{b}_{H+1})\}$.
@@
Given a vector of observation $o\in\mathbb{R}^{N\times d}$, the neural network outputs a vector $g_{\theta}(o)=(g_{\theta}(o_{1}),\dots,g_{\theta}(o_{N}))\in\mathbb{R}^{N}$. Then, the action $a=\operatorname*{argmax}_{i\in\llbracket 1,N\rrbracket}g_{\theta}(o_{i})$ is returned by $\pi_{\theta}$.
Figure 2: Neuro-LS policy
@@
The neural network policy $\pi_{\theta}$ is characterized by a set of parameters denoted as $\theta$. The optimization goal is to maximize the estimated score $\bar{F}(\pi_{\theta},{\cal NK}(N,K),H)$. This poses a stochastic black-box optimization problem within the real-valued search space $\mathbb{R}^{\mid\theta\mid}$. To tackle this problem, we propose to use the covariance matrix adaptation evolution strategy (CMA-ES) hansen2001completely , which stands out as one of the most powerful evolutionary algorithms for addressing such black box optimization problems muller2018challenges .

### v2

@@ Abstract
This paper explores a novel approach aimed at overcoming existing challenges in the realm of local search algorithms. Our aim is to improve the decision process that takes place within a local search algorithm so as to make the best possible transitions in the neighborhood at each iteration. To improve this process, we propose to use a neural network that has the same input information as conventional local search algorithms. In this paper, which is an extension of the work presented at EvoCOP2024, we investigate different ways of representing this information so as to make the algorithm as efficient as possible but also robust to monotonic transformations of the problem objective function. To assess the efficiency of this approach, we develop an experimental setup centered around NK landscape problems, offering the flexibility to adjust problem size and ruggedness. This approach offers a promising avenue for the emergence of new local search algorithms and the improvement of their problem-solving capabilities for black-box problems. The last version of this article is published in the journal SN Computer Science (Springer).
@@ 1 Introduction
Fitness landscape analysis malan2021 provides the optimization and evolutionary computation community with measurable characteristics (e.g, ruggedness or presence of local optima networks) and practical tools (e.g., visualization frameworks) to examine search landscapes. It allows for the assessment of problem characteristics and the evaluation of the performance of search algorithms. In the context of combinatorial fitness landscapes, these are represented as graphs defined by a discrete search space and a neighborhood relation. A fundamental challenge lies in developing a search algorithm capable of navigating a fitness landscape to reach the highest possible fitness value.
@@
In various works, different machine learning approaches have been used either to consider offline adjustment, selection of parameters, or online control of the search process using reinforcement learning (RL) techniques (see MamaghanMMKT22 for a recent survey). The use of neural networks (NNs) in solving combinatorial optimization problems has been studied for decades TrafalisK09 , starting with the early work of Hopfield Hopfield82 . Recent applications of Graph Neural Networks (GNN) in the context of combinatorial optimization have been proposed to reach optimal solutions or to assist the solving algorithm in proving the optimality of a given solution (see CappartCK00V23 for a recent survey).
@@
In the traveling salesman problem, a GNN can be used to predict the regret [^3] In machine learning, regret measures the difference between the actual cumulative reward obtained by a learning process and the cumulative reward of an optimal strategy. It quantifies how much worse the algorithm performs compared to the best possible choice. Hence, lower regret indicates better performance. associated with adding each edge to the solution to improve the computation of the fitness function of the LS algorithm HudsonLMP22 . Note that NNs are classically used in surrogate model-based optimization WillmesBJS03 . In SantanaLV23 , the authors introduce a GNN into a hybrid genetic search process to solve the vehicle routing problem. The GNN is used to predict the efficiency of search operators and to select them optimally in the solving process. A deep Q-learning approach has recently been proposed to manage the different stages of an LS-based metaheuristic to solve routing and job-shop problems FalknerTBS22 . High-level solving policies can often be managed by reinforcement learning in LS processes VeerapenHS13 . In this paper, our purpose is different, as we focus on building simple search heuristics for black-box problems, rather than scheduling specific operators or parameters. In particular, we assume that the learning process cannot be based on the immediate rewards that are used in RL. This motivates our choice of neuro-evolution. Note that this type of approach has recently been used to discover new genetic and evolutionary algorithms to solve continuous black-box optimization problems lange2023discoveringGen ; lange2023discoveringEvo . In this work, we propose to use it to discover new local search algorithms for combinatorial optimization problems.
@@
Objective of the paper This study explores the potential for emerging search strategies to overcome existing challenges. The objective is to change how information is leveraged while retaining simple and generic search components. Considering a basic hill climber algorithm to achieve a baseline search process for solving black-box binary problem instances, our aim is to benefit from machine learning techniques to get new local search heuristics that will be built from basic search information instead of choosing a priori a predefined move heuristic (e.g., always select the best neighbor). Hence, our goal is to provide a NN with the same information as a basic LS and, after training, to use the NN as the basic move component of a simple LS process. In this work, we study different ways of representing this information, from the most basic possible representation to a more complex one that exploits an ordering relation in the neighborhood. To evaluate the efficiency of our approach, we define an experimental setup based on NK landscape problems kauffman1989 , which allows us to describe a fitness landscape whose problem size and ruggedness (determining the number of local minima) can be adjusted as parameters. Finally, to check the robustness and generalizability of our learned strategies, we propose to evaluate them on new instances of another type of pseudo-Boolean problem, namely the quadratic unconstrained binary optimization (QUBO) problem.
@@ 2.2.2 Observations
In this work, we propose to study different observation functions that will be used by local search strategies to make decisions. Given $s\in\mathcal{S}$, with $s=(x,\mathcal{N}(x))$, with $x$ the current solution encountered by the local search and $\mathcal{N}(x)=\{flip_{i}(x)\mid i\in\llbracket 1,N\rrbracket\}$ its one-flip neighborhood, we define four observation functions from $o^{k}:\mathcal{S}$ to $\mathbb{R}^{N\times d}$ using superscript $k$ to distinguish between them. The notation $o^{k}(x)_{i}$ is used to denote the $i^{th}$ line of $o^{k}(x)$ (in case of $d=1$ this line if of course the $i^{th}$ value of the corresponding 1D vector).
@@
    The first observation function abstracts the state as the variation of the fitness function for each possible flip of the values of $x$ ($d$ is set to 1).
    $o^{1}(x)=\begin{pmatrix}\Delta_{1}(x)\\ \Delta_{2}(x)\\ \vdots\\ \Delta_{N}(x)\\ \end{pmatrix},$
@@
    $o^{2}(x)=\begin{pmatrix}f_{{\rm obj}}(x)&f_{{\rm obj}}(flip_{1}(x))\\ f_{{\rm obj}}(x)&f_{{\rm obj}}(flip_{2}(x))\\ \vdots&\vdots\\ f_{{\rm obj}}(x)&f_{{\rm obj}}(flip_{N}(x))\\ \end{pmatrix},$
    Let us note that the same value $f_{{\rm obj}}(x)$ is duplicated in each line of $o^{2}(x)$ since the NN (see Section 3) is applied to each line $o(x)_{i}$.
@@
    The third observation function $o^{3}:\mathcal{S}\rightarrow\mathbb{R}^{{N}\times{d}}$, with $d=1$ is based on a ranking of the solutions associated with positive, negative or null variations of the fitness function for each possible flip of the values of $x$. We define $N^{+}(x)$ and $N^{-}(x)$ the numbers of solutions in $\mathcal{N}(x)$ with respectively strictly greater and lower fitness than $x$.
    Hence, $o^{3}(x)$ is vector of size $N$, such that for $i=1,\dots,N$:
    • $o^{3}(x)_{i}=rank^{+}(\Delta_{i}(x))/N^{+}(x)$, if $\Delta_{i}(x)>0$;
    • $o^{3}(x)_{i}=rank^{-}(\Delta_{i}(x))/N^{-}(x)$, if $\Delta_{i}(x)<0$;
    • $o^{3}(x)_{i}=0$, if $\Delta_{i}(x)=0$.
@@
    Example 2 Let us consider the following situation with $x\in S$ and $N=6$ : $\Delta_{1}(x)=1,\Delta_{2}(x)=4,\Delta_{3}(x)=-2,\Delta_{4}(x)=-5,\Delta_{5}(x)=0,\Delta_{6}(x)=-7$. We have thus $rank^{+}(\Delta_{1}(x))=1,rank^{+}(\Delta_{2}(x))=2,rank^{-}(\Delta_{3}(x))=-1,rank^{-}(\Delta_{4}(x))=-2,rank^{-}(\Delta_{6}(x))=-3$. Hence, the corresponding observations are $o^{3}(\Delta_{1}(x))=\frac{1}{2},o^{3}(\Delta_{2}(x))=1,o^{3}(\Delta_{3}(x))=-\frac{1}{3},o^{3}(\Delta_{4}(x))=-\frac{2}{3},o^{3}(\Delta_{5}(x))=0,o^{3}(\Delta_{6}(x))=-1$.
    This rank transformation always returns values in the range $[-1,1]$ and keeps the information that a flip move can deteriorate or improve the current fitness solution. An interesting property of this transformation, is that it is independent of a change in the scale of the objective function. $o^{3}(x)$ remains unchanged even if $f_{{\rm obj}}$ is multiplied by a strictly positive real value $\lambda$.
@@
    The fourth observation function $o^{4}:\mathcal{S}\rightarrow\mathbb{R}^{N\times d}$, with $d=2$ uses the same ranking transformation as $o^{3}$, in addition to a z-score transformation of the $\Delta_{i}(x)$ values. We define the z-score of $\Delta_{i}(x)$ as $Z(\Delta_{i}(x))=\frac{\Delta_{i}(x)-\mu(x)}{\sigma(x)}$, with $\mu(x)$ and $\sigma(x)$ are the mean and standard deviation, respectively, of the set of $\Delta_{i}(x)$ values for $i=1,\dots,N$. This z-score indicates how many standard deviations a given fitness variation is above or below the mean of possible fitness variations. Thus,
    $o^{4}(x)=\begin{pmatrix}o^{3}(x)_{1}&Z(\Delta_{1}(x))\\ o^{3}(x)_{2}&Z(\Delta_{2}(x))\\ \vdots&\vdots\\ o^{3}(x)_{N}&Z(\Delta_{N}(x))\\ \end{pmatrix}.$
@@ 2.2.3 Deterministic policy and actions
Following a reinforcement learning-based description, a local search process can be encoded by a policy $\pi:\mathbb{R}^{N\times d}\to{\cal A}$ where ${\cal A}$ is a set of actions. Here we consider deterministic policies, i.e. for each $o(x)\in\mathbb{R}^{N\times d}$, there exists one and only one action $a\in{\cal A}$, such that $\pi(o(x))=a$. Note that the policy can be parameterized by a parameter vector $\theta$. The set of all policies is $\Pi=\{\pi_{\theta}\mid\theta\in\Theta\}$ where $\Theta$ is the parameter space.
@@ 2.2.4 Transitions and trajectory
Example 1. In order to illustrate our framework, let us consider a basic hill climber (HC) that uses a simple best-improve strategy using objective function $f_{{\rm obj}}$ as fitness function. In the following, we only consider LS processes that do not involve memory that records past decisions. Hence, a state of a HC is fully described by a current solution $x\in{\cal X}$ and its neighborhood $\mathcal{N}(x)$. An observation will abstract the state as the variation of the fitness function for each possible flip of the values of $x$, $o^{2}(x)=(\Delta_{1}(x),\dots,\Delta_{N}(x))\in\mathbb{R}^{N}$. In a best-improve HC process, the search stops when a local optimum is reached and no improving move can be performed. We consider ${\cal A}=\{flip_{i}\mid i\in\llbracket 1,N\rrbracket\}$. Then, we define the policy $\pi_{HC}(o^{1}(x))=\operatorname*{argmax}_{a\in{\cal A}}(f_{{\rm obj}}(\delta(x,a))-f_{{\rm obj}}(x))$. Note that when a local optimum is reached, the search gets stuck.
@@ 2.3 Policy Learning for a Set of Instances
Figure 1: Global View of the Process.
@@ 3.1 Neural network local search policy
We introduce a deterministic LS policy $\pi_{\theta}:\mathbb{R}^{N\times d}\rightarrow\mathcal{A}$, called Neuro-LS, which uses a neural network $g_{\theta}$, parametrized by a vector of real numbers $\theta$. When in a state $x$, this neural network takes as input an observation matrices $o(x)\in\mathbb{R}^{N\times d}$ and gives as output a vector $g_{\theta}(o(x))$ of size $N$ whose component $g_{\theta}(o(x))_{i}$ corresponds to a preference score associated to each observation $o_{i}$. Then, the action $a_{i}$ corresponding to the highest score $g_{\theta}(o)_{i}$ is selected.
@@
A desirable property of this neural network policy is to be permutation equivariant with respect to the input vector of observations, which is a property generally entailed by a local search algorithm, in order to make its behavior consistent for solving any type of instance. Formally, an LS algorithm is said to be invariant to permutations in the observations if for any permutation $\sigma$ on $\llbracket 1,N\rrbracket$, we have $a_{\sigma(i)}=\pi_{\theta}(o(x)_{\sigma(1)},o(x)_{\sigma(2)},\dots,o(x)_{\sigma(N)})$. As an example, the basic hill climber HC defined with Example 1 in Section 2.2 has this property.
@@
where $L_{h}:\mathbb{R}^{n_{h-1}}\rightarrow\mathbb{R}^{n_{h}}$ is an affine linear map defined by $L_{h}({x})={W}_{h}\cdot{x}+{b}_{h}$ for given $n_{h}\times n_{h-1}$ dimensional weight matrix ${W}_{h}$ (with coefficients $\{w^{h}_{k,l}\}_{1\leq k\leq n_{h}\atop 1\leq l\leq n_{h-1}}$), $n_{h}$ dimensional bias vector ${b}_{h}$ (with coefficients $\{b^{h}_{k}\}_{1\leq k\leq n_{h}}$) and $\sigma:\mathbb{R}^{n_{h}}\rightarrow]-1,1[^{n_{h}}$444]1,1[]-1,1[] - 1 , 1 [ is used to denote the open interval included in $\mathbb{R}$, while $\llbracket 1,N\rrbracket$ in an integer values interval. the element-wise nonlinear activation map defined by $\sigma({z}):=(tanh(z_{1}),...,tanh(z_{n_{h}}))^{\intercal}$. We denote by $\theta$, the set of all weight matrices and bias vector of the neural network $g_{\theta}$ : $\theta:=\{({W}_{1},{b}_{1}),({W}_{2},{b}_{2}),\dots,({W}_{{H+1}},{b}_{H+1})\}$.
@@
Given a vector of observation $o(x)\in\mathbb{R}^{N\times d}$, the neural network outputs a vector $g_{\theta}(o(x))=(g_{\theta}(o(x)_{1}),\dots,g_{\theta}(o(x)_{N}))\in\mathbb{R}^{N}$. Then, the action $a=\operatorname*{argmax}_{i\in\llbracket 1,N\rrbracket}g_{\theta}(o(x)_{i})$ is returned by $\pi_{\theta}$.
Figure 2: Neuro-LS policy .
@@
The neural network policy $\pi_{\theta}$ is characterized by a set of parameters denoted as $\theta$. The optimization goal is to maximize the estimated score $\bar{F}(\pi_{\theta},{\cal NK}(N,K),H)$. This poses a stochastic black-box optimization problem within the real-valued search space $\mathbb{R}^{\mid\theta\mid}$. To tackle this problem, we propose to use the covariance matrix adaptation evolution strategy (CMA-ES) hansen2001completely , which stands out as one of the most powerful evolutionary algorithms for addressing such black-box optimization problems muller2018challenges .